OpenStudy (anonymous):

Weights of male mountain lions follow the normal distribution with a median of 150 lb and an interquartile range of 8.2 lb, Find the 75th percentile of the weights.

OpenStudy (perl):

median?

OpenStudy (perl):

ok this is a trick question. in a normal distribution the median=mean=mode. so we have a normal distribution with a mean of 150 lb. the interquartile range is 75th percentile - 25th percentile

OpenStudy (perl):

Q3 = 75th percentile, Q1 = 25th percentile We are given the interquartile range is Q3 - Q1 = 8.2 , half of this is 4.1 , so Q3 - Q2 = 4.1, where Q2 is the mean and Q3 is the 75th percentile Q3 - 150 = 4.1 Q3 = 150 + 4.1 = 154.1